OpenStudy (lukecrayonz):

Find the area under the graph of f over [-2,2] f(x)={4-x^2, if x<0 4, if x>0

OpenStudy (lukecrayonz):

Sorry, \[4 if x \ge 0\]

OpenStudy (lukecrayonz):

I got 9.3, not sure where I went wrong.

OpenStudy (btaylor):

It looks as if you will need to set up 2 separate integrals. First, integrate the top function (for x<0) on [-2,0]. --> 16/3 Then, add it to the integral of the bottom part on [0,2]. --> 8 Your exact answer is 40/3, or 13.333333333333...

OpenStudy (anonymous):

YES. 40/3

OpenStudy (lukecrayonz):

Okay :O So then for the area bounded by y=x^2 and y=4

OpenStudy (lukecrayonz):

Intersects are (-2,4) and (2,4) so then my interval is -[2,2]

OpenStudy (lukecrayonz):

\[\int\limits_{-2}^{2}x^2dx-\int\limits_{-2}^{2}4dx\]